# DistALANER: Distantly Supervised Active Learning Augmented Named Entity Recognition in the Open Source Software Ecosystem

### v3

@@ 12 Error analysis
NER can help extract entities like library names, function names, software versions, bug IDs, or even contributor names from this sea of information. However, it’s vital to ensure the NER system’s accuracy, and this is where error analysis comes in. In this section, we analyze the incorrect predictions and group them into the following types.
  • Ambiguity errors: These occur when a token could be identified as multiple entity types or when it’s unclear whether a token is an entity. For example, "Apple" could refer to the OS or an organization in a bug. Enhancing training data to better cover such ambiguous instances or using more advanced models that are adept at handling ambiguities can help.
  • Out-of-vocabulary (OOV) errors: These occur when the system encounters tokens that were not included in the training data. In the rapidly evolving open-source domain, new software, libraries, and tools frequently appear, resulting in unfamiliar tokens. Character-based or subword-based NER can be used to handle words not seen during training.
  • Boundary detection errors: These occur when the system inaccurately determines the start and end of an entity. For example, the system might identify "Windows NT3.‘..’" as separate entities instead of a single entity referring to a specific version of the Windows version. Solutions can include using techniques like BIO tagging (Begin, Inside, Outside) or increasing the range of training examples with varying entity lengths.
  • Incorrect Entity Type Errors: These occur when the NER system correctly identifies an entity but assigns it the wrong type. For instance, mistaking a software version for a date. Training models on a diverse set of examples and refining entity type definitions can mitigate this issue.
  • Cohesion errors: These are related to the semantics of software domain language. For instance, "sudo, apt, update" should be recognized as individual instead of whole "sudo apt update". This problem can be mitigated by employing sequence tagging models and enhancing the training set with more examples of such entities.
  • Errors due to homonyms: Some words may have different meanings depending on the context. For example, "Java" can refer to a programming language, an island, or a type of coffee. Models need to be trained to understand the context to disambiguate the intended meaning.
  • Multi Annotator disagreement: In the domain of open source software, disagreements among multiple expert annotators can frequently arise due to the inherent complexity and variety of software artifacts. These disagreements can be a result of different interpretations of code, differing understanding of software architecture, or varying familiarity with the programming languages and software systems in question. This heterogeneity in expert opinions can pose significant challenges in tasks such as bug labeling, feature identification, or documentation generation.
@@
13 Conclusion
14 Limitation
While the distant supervision improves NER performance it relies on an existing knowledge base to automatically annotate data. If this knowledge base contains errors or is incomplete, this can directly affect the quality of the supervision and thus the performance of the model. Further distantly supervised models may struggle with multi-token entities (e.g., software names with versions like “Ubuntu 20.04 LTS”). Boundary detection for these entities can be problematic, leading to incomplete or incorrect entity recognition. Finally, errors in the NER stage can propagate to the relation extraction stage. If an entity is not correctly recognized, any relation involving that entity will inherently be incorrect.
15 Ethical consideration
The information we gather in our dataset is free from harmful or offensive materials. We take serious measures to anonymize and handle any personal or sensitive data with the highest level of confidentiality. Protection of participants’ privacy is our priority and we consistently ensure to acquire their informed consent when collecting, annotating, and analyzing data. We provide equal incentives to all annotators for their efforts toward the annotation task.
Appendix A Prompt
@@
Appendix B Regex details
@@
Appendix C Interannotator agreement
@@
Appendix D Hyperparameter settings

### v4

@@ 12 Error analysis
In this section, we analyze the incorrect predictions and group them into the following types.
Ambiguity errors: Occur when a token’s meaning is unclear, such as "Apple" referring to either the OS or an organization. Solutions include enriching training data for ambiguous cases and utilizing advanced models proficient in ambiguity resolution.
Out-of-Vocabulary (OOV) errors: This arises with tokens not in training data, common in the evolving open-source field. Employing character-based or subword-based NER techniques can address OOV issues.
Boundary detection errors: This happens when entity boundaries are wrongly identified, e.g., mistaking "Windows NT3.‘.." for separate entities. BIO tagging or expanding training examples with varied entity lengths can help.
Incorrect entity type errors: When entities are correctly identified but misclassified, such as a software version labeled as a date. Broadening the diversity of training examples and refining entity definitions can reduce these errors.
Cohesion errors: Related to software domain semantics, e.g., "sudo, apt, update" should be separate entities. This can be addressed by using sequence tagging models and enriching the training dataset.
Errors due to homonyms: Words with multiple meanings like "Java" need context-aware models to resolve ambiguities based on usage.
Multi-annotator disagreement: Occurs when expert annotators have differing opinions due to the complexity and diversity of software artifacts. Addressing this requires acknowledging and accommodating the range of expert interpretations.
13 Conclusion
Appendix 0.A Error Analysis
@@
Appendix 0.B Prompt
@@
Appendix 0.C Regex details
@@
Appendix 0.D Interannotator agreement
@@
Appendix 0.E Hyperparameter settings